Calculate the change in Gibbs free energy for each of the following sets of ΔH∘rxn, ΔS∘rxn, and T, assuming that all reactants and products are in their standard states.

ΔH∘rxn = −90.kJ, ΔS∘rxn = −150J/K, T = 301K

ΔH∘rxn = 90.kJ, ΔS∘rxn = −150J/K, T = 301K

ΔH∘rxn = −90.kJ, ΔS∘rxn = −150J/K, T = 856K

ΔH∘rxn = −90.kJ, ΔS∘rxn = 150J/K, T = 401K
Answer #1

1)

ΔHo = -90.0 KJ

ΔSo = -150 J/K

= -0.15 KJ/K

T = 301 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -90.0 - 301.0 * -0.15

ΔGo = -44.85 KJ

Answer: -44.85 KJ

2)

ΔHo = 90.0 KJ

ΔSo = -150 J/K

= -0.15 KJ/K

T = 301 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = 90.0 - 301.0 * -0.15

ΔGo = 135.15 KJ

Answer: 135.15 KJ

3)

ΔHo = -90.0 KJ

ΔSo = -150 J/K

= -0.15 KJ/K

T = 856 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -90.0 - 856.0 * -0.15

ΔGo = 38.4 KJ

Answer: 38.4 KJ

4)

ΔHo = -90.0 KJ

ΔSo = 150 J/K

= 0.15 KJ/K

T = 401 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -90.0 - 401.0 * 0.15

ΔGo = -150.15 KJ

Answer: -150.15 KJ